OpenStudy (anonymous):

Wake Up coffee shop offers two different sizes of coffee. The small coffee costs $2 and the grande coffee costs $3. It is your turn to buy coffee for the office and the bill came to $18. This scenario can be modeled by the equation 2s + 3g = 18. Which is not a possible combination of small and grande coffees? Choose one answer. a. 9 small and 1 grande b. 0 small and 6 grande c. 3 small and 4 grande d. 6 small and 2 grande is this a?

OpenStudy (anonymous):

A. is not possible because 2(9) + 3(1) = 21. B. 2(0) + 3(6) = 18 C. 2(3) + 3(4) = 18 D. 2(6) + 3(2) =18 All of the answer choices equal 18 except A.